OpenStudy (anonymous):

Need help..... -32=-7n-12+3n

OpenStudy (anonymous):

Well the first thing you should do is combine like terms. On the right side of the equation, you have -7n and 3n. If you add those together, you get 4n

OpenStudy (anonymous):

-4n

OpenStudy (anonymous):

Sorry about that

OpenStudy (anonymous):

So you have -32=-4n-12

OpenStudy (anonymous):

Then you can add 12 to the other side, to isolate the term with n on the other side, as well as to combine more like terms

OpenStudy (anonymous):

So you get -20=-4n

OpenStudy (anonymous):

Then if you divide both sides by -4, you get that n=5

OpenStudy (anonymous):

This is correct because if you substitute 5 into the original equation, you get

OpenStudy (anonymous):

Thank You!!!!

OpenStudy (anonymous):

-32=-7(5)-12+3(5)

OpenStudy (anonymous):

Which comes out to -32=-35-12+15, which is -32=-32, which is correct :)

OpenStudy (anonymous):

No problem!!